OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (–2, –3) with a slope of –6. (5 points)

OpenStudy (anonymous):

helllllllllllllllllppp

OpenStudy (anonymous):

\[y – y_{1} = m(x – x_{1})\]\[x_{1} = -2; y_{1} = -3; m=-5\] Hope that helps.

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

oops, m=-6... the problem is worth 5

OpenStudy (anonymous):

ya sorry about that m=-6. just keeping you on your toes. ;)